OpenStudy (anonymous):

Given the linear equation x – y = 5, find the slope of its graph. (Points : 5) –5 –1 1 5

hartnn (hartnn):

can you isolate y from that equation ?

OpenStudy (anonymous):

slope of graph, or gradient is given by, y = mx +c,

hartnn (hartnn):

bring y-x = 5 in the form of y= mx+c then m= slope